40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 22     Entry time: Sun Oct 28 03:03:42 2007
Author: rana 
Type: Configuration 
Category: IOO 
Subject: Three Way Excitement 
We've been trying to measure the MC mirror internal mode frequencies so that we can measure
their absorption before and after drag wiping.


It looked nearly impossible to see these modes as driven by their thermal excitation level;
we're looking at the "MC_F" or 'servo' output directly on the MC servo board.

Today, I set up a band limited noise drive into the 'Fast POS' inputs of the 3 MC coil
driver boards (turns out you can do this with either the old HP or the SR785).

Frequencies:
MC1     28.21625 kHz
MC2     28.036   kHz
MC3     28.21637 kHz

I don't really have this kind of absolute accuracy. These are just numbers read off of the SR785.

The other side of the setup is that the same "MC_F" signal is going into the SR830 Lock-In which
is set to 'lock-in' at 27.8 kHz. The resulting demodulated 'R" signal (magnitude) is going into
our MC_AO channel (110B ADC).

As you can see from the above table, MC1 and MC3 are astonishingly and annoyingly very close in
frequency. I identified mirrors with peaks by driving one at a time and measuring on the spectrum
analyzer. I repeated it several times to make sure I wasn't fooling myself; it seems like they
are really very close
but distinct peaks. I really wish we had chipped one of these mirrors
before installing them.



Because of the closeness of these drumhead modes, we will have to measure the absorption by making long
measurements of this channel.
ELOG V3.1.3-